### v1

@@ Abstract
This paper presents the winning solution of task 1 and the third-placed solution of task 3 of the BraTS challenge. The use of automated tools in clinical practice has increased due to the development of more and more sophisticated and reliable algorithms. However, achieving clinical standards and developing tools for real-life scenarios is a major challenge. To this end, BraTS has organised tasks to find the most advanced solutions for specific purposes. In this paper, we propose the use of synthetic data to train state-of-the-art frameworks in order to improve the segmentation of adult gliomas in a post-treatment scenario, and the segmentation of meningioma for radiotherapy planning. Our results suggest that the use of synthetic data leads to more robust algorithms, although the synthetic data generation pipeline is not directly suited to the meningioma task. The code for these tasks is available at https://github.com/ShadowTwin41/BraTS_2023_2024_solutions.
@@ 2.2 Data augmentation
The same approach as in [3] was used to train two different GliGANs for both Task 1 and Task 3 in order to insert synthetic tumours into the healthy part of the scans. This strategy aims to increase the variability of tumours and reduce the class imbalance between healthy and unhealthy tissue. A random label generator based on [2] was also trained to generate random labels in both tasks and to be used in the trained GliGANs. We randomly choose whether to use a real label or a generated label for inference.
@@
It is important to note that each GAN was trained individually, i.e. no data from one dataset was used to train the GAN of the other dataset. Both tasks were developed independently of each other.
@@ 3 Results
Table 1: Validation set results computed by the validation platform. The "All" is defined as $rGN_{g}RN_{g}rGS_{g}RS_{g}rGM_{g}RM_{g}$. The "All real" is composed by all models only trained with real data. The "All fake" is composed by all models trained with real and synthetic data.
@@ 4 Discussion and Conclusion
The results of the validation set for tasks 1 and 3 are presented in tables 1 and 2.
@@
Regarding task 1, from table 1 it can be seen that it is still room for improvement, specially of the RC region which had the best DSC around 0.7. Similarly to BraTs 2023 Task 1 solution [3], the ensemble of several models shows a very good performance compared to each individual model, as the number of false positives (FP) is reduced without significantly increasing the number of FN. The models trained with synthetic data, when tested individually, have a lower number of FN, however a considerably higher number of FP which heavily penalises the results. The random generation of synthetic tumours anywhere in the brain might therefore potentially make the training process for the segmentation very unstable and the model very prone to FP, as too many non-existent tumours are predicted, specially the Swin UNETR. The models trained with MexNeXt have less FP than the others, but more FN. Therefore, it was determined that the ensemble of all models would have the greatest potential for the test phase as it was balanced in terms of FP and FN. We decided on a threshold of 50 for WT and 50 for RC in the post-processing step. Using larger thresholds produced worst results, meaning that our final solution did not produce many small FP. With this solution, we have achieved the DSC and HD95 specified in the first entry of the table 1 in the validation phase.
@@
For task 3 a semi-lesion-wise metric was used to evaluate the predictions, i.e., the evaluation metric does not penalise the FP as these can be easily identified and removed by a treating radiation oncologist. However, the FN are still penalised. In this case, the use of a threshold does not improve much the results, as can be seen in Table 2. It can even produce worse results by removing predictions that are true positives, creating FN. Unlike in Task 1, there is no good balance between the number of FP and FN. The use synthetic data to train the MedNeXt framework had a positive effect on the results. The number of FN is the lowest among all the solutions, which had a direct effect on the lesion-wise DSC. As mentioned in the section "2.2 Data augmentation", the pipeline for generating synthetic data was not fully adapted to this dataset. However, the randomness of the locations where the synthetic tumour was located made the network more sensible and robust, avoiding FN. The most successful model tested was then the MexNeXt trained with real and synthetic data and no threshold, attaining a DSC of 0.8214 and HD95 of 24.64 in the validation phase. Furthermore, this version of the model was able to achieve the lowest number of FN of the ones tested.
@@
For future work, it would be important to further improve the GliGAN pipeline to select better locations for the placement of the synthetic tumour and to generate more realistic cases.

### v2

@@ Abstract
This paper presents the winning solution of task 1 and the third-placed solution of task 3 of the BraTS challenge. The use of automated tools in clinical practice has increased due to the development of more and more sophisticated and reliable algorithms. However, achieving clinical standards and developing tools for real-life scenarios is a major challenge. To this end, BraTS has organised tasks to find the most advanced solutions for specific purposes. In this paper, we propose the use of synthetic data to train state-of-the-art frameworks in order to improve the segmentation of adult gliomas in a post-treatment scenario, and the segmentation of meningioma for radiotherapy planning. Our results suggest that the use of synthetic data leads to more robust algorithms, although the synthetic data generation pipeline is not directly suited to the meningioma task. In task 1, we achieved a DSC of 0.7900, 0.8076, 0.7760, 0.8926, 0.7874, 0.8938 and a HD95 of 35.63, 30.35, 44.58, 16.87, 38.19, 17.95 for ET, NETC, RC, SNFH, TC and WT, respectively and, in task 3, we achieved a DSC of 0.801 and HD95 of 38.26, in the testing phase. The code for these tasks is available at https://github.com/ShadowTwin41/BraTS_2023_2024_solutions.
@@ 2.2 Data augmentation
The same approach as in [3] was used to train two different GliGANs for both Task 1 and Task 3 in order to insert synthetic tumours into the healthy part of the scans, as shown in Figure 1. This strategy aims to increase the variability of tumours and reduce the class imbalance between healthy and unhealthy tissue. A random label generator based on [2] was also trained to generate random labels in both tasks and to be used in the trained GliGANs. We randomly choose whether to use a real label or a generated label for inference.
Figure 1: The training pipeline of the GliGAN. The noise scan (z) and the label (y) are concatenated and fed into the generator (G). The discriminator (D) assesses the realism of a real scan and the reconstruction. Figure taken from [3].
@@
It is important to note that each GAN was trained individually, i.e. no data from one dataset was used to train the GAN of the other dataset. Both tasks were developed independently of each other. To ensure the realism of the synthetic tumours, visual inspections were performed at several checkpoints. The loss graphs were also observed to ensure that the training of the GAN did not collapse. We observed very stable training in all situations.
@@ 3 Results
Table 1: Validation set results for Task 1 computed by the validation platform. The "All" is defined as $rGN_{g}RN_{g}rGS_{g}RS_{g}rGM_{g}RM_{g}$. The "All real" is composed by all models only trained with real data. The "All fake" is composed by all models trained with real and synthetic data.
@@
Table 3: Test set results for Task 1 computed by the challenge organisers.
Region DSC (Mean) DSC (Std) HD95 (Mean) HD95 (Std)
ET 0.7900 0.2958 35.63 89.86
NETC 0.8076 0.3207 30.35 90.15
RC 0.7760 0.3263 44.58 106.78
SNFH 0.8926 0.1591 16.87 47.48
TC 0.7874 0.2883 38.19 90.61
WT 0.8938 0.1634 17.95 49.36
Table 4: Test set results for Task 3 computed by the challenge organisers.
Region DSC (Mean) DSC (Std) HD95 (Mean) HD95 (Std)
GTV 0.801 0.221 38.26 139.45
@@ 4 Discussion and Conclusion
The results of the validation set for tasks 1 and 3 are presented in tables 1 and 2. The test results of task 1 and 3 provided by the organisers are presented in tables 3 and 4.
@@
Regarding task 1, from table 1 it can be seen that it is still room for improvement, specially of the RC region which had the best DSC around 0.7. Similarly to BraTs 2023 Task 1 solution [3], the ensemble of several models shows a very good performance compared to each individual model, as the number of false positives (FP) is reduced without significantly increasing the number of FN. The models trained with synthetic data, when tested individually, have a lower number of FN, however a considerably higher number of FP which heavily penalises the results. The random generation of synthetic tumours anywhere in the brain might therefore potentially make the training process for the segmentation very unstable and the model very prone to FP, as too many non-existent tumours are predicted, specially the Swin UNETR. The models trained with MexNeXt have less FP than the others, but more FN. Therefore, it was determined that the ensemble of all models would have the greatest potential for the test phase as it was balanced in terms of FP and FN. We decided on a threshold of 50 for WT and 50 for RC in the post-processing step. Using larger thresholds produced worst results, meaning that our final solution did not produce many small FP. With this solution, we have achieved the DSC and HD95 specified in the first entry of the table 1 in the validation phase and the best ranking in the testing phase. The DSC and HD95 of the testing phase are presented in table 3.
@@
For task 3 a semi-lesion-wise metric was used to evaluate the predictions, i.e., the evaluation metric does not penalise the FP as these can be easily identified and removed by a treating radiation oncologist. However, the FN are still penalised. In this case, the use of a threshold does not improve much the results, as can be seen in Table 2. It can even produce worse results by removing predictions that are true positives, creating FN. Unlike in Task 1, there is no good balance between the number of FP and FN. The use synthetic data to train the MedNeXt framework had a positive effect on the results. The number of FN is the lowest among all the solutions, which had a direct effect on the lesion-wise DSC. As mentioned in the section "2.2 Data augmentation", the pipeline for generating synthetic data was not fully adapted to this dataset. However, the randomness of the locations where the synthetic tumour was located made the network more sensible and robust, avoiding FN. The most successful model tested was then the MexNeXt trained with real and synthetic data and no threshold, attaining a DSC of 0.8214 and HD95 of 24.64 in the validation phase. Furthermore, this version of the model was able to achieve the lowest number of FN of the ones tested. Therefore, this was the solution submitted to the testing phase, archiving the results presented in table 4.
@@
For future work, it would be important to further improve the GliGAN pipeline to select better locations for the placement of the synthetic tumour and to generate more realistic cases. Statistical similarity assessments should also be used to evaluate the quality of the synthetic tumours, rather than just a visual inspection.